Solve the system of equations
x+8y=9
x=6-8y

Answer by ewatrrr(24785)


Hi,
x+8y=9
x=6-8y
x+8y=9
x+8y=6
x+8y=9
-x-8y=-6 Multiplying the second eqation by -1 and adding them together
0 = 3
As this is not true, there are no solutions for this system